### v2

# What is the Plausible Memory Representation?

Abstract
Recent work in computational psycholinguistics has revealed intriguing parallels between attention mechanisms and human memory retrieval, focusing primarily on vanilla Transformers that operate on token-level representations. However, computational psycholinguistic research has also established that syntactic structures provide compelling explanations for human sentence processing that token-level factors cannot fully account for. In this paper, we investigate whether the attention mechanism of Transformer Grammar (TG), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE) as a linking hypothesis between models and humans. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general memory retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.
If Attention Serves as a Cognitive Model of Human Memory Retrieval,
What is the Plausible Memory Representation?
Ryo Yoshida Shinnosuke Isono Kohei Kajikawa Taiga Someya Yushi Sugimoto Yohei Oseki The University of Tokyo NINJAL The University of Osaka {yoshiryo0617,taiga98-0809,oseki}@g.ecc.u-tokyo.ac.jp {s-isono,kohei-kajikawa}@ninjal.ac.jp sugimoto.yushi.hmt@osaka-u.ac.jp
1 Introduction
Whether language models (LMs) developed in natural language processing (NLP) are plausible as cognitive models of human sentence processing is a central question in computational psycholinguistics. Over the past two decades, this question has been primarily addressed from the perspective of expectation-based theories—one of the two major classes of human sentence processing theory—examining whether LMs’ next-token prediction can serve as a model of human predictive processing (Hale, 2001; Levy, 2008; Wilcox et al., 2020; Merkx and Frank, 2021; inter alia).
The recent success of Transformers Vaswani et al. (2017) in NLP has unexpectedly opened a new avenue of investigation from the perspective of memory-based theories, the other major class of sentence processing theory. Researchers have proposed that the attention mechanism, despite its engineering origins, can implement a human memory retrieval theory known as cue-based retrieval Van Dyke and Lewis (2003). Recent studies have revealed intriguing parallels between the weighted reference patterns exhibited by the attention mechanism and the elements that humans may retrieve during online sentence comprehension (Ryu and Lewis, 2021; Oh and Schuler, 2022; Timkey and Linzen, 2023).
Computational psycholinguistics has also established that human sentence processing cannot be fully explained by token-level factors; rather, syntactic structures have provided compelling explanations for it. For instance, next-token prediction from LMs that explicitly incorporate syntactic structure building demonstrates superior performance in accounting for human brain activity compared to vanilla RNNs and Transformers (Hale et al., 2018; Wolfman et al., 2024); the number of syntactic nodes hypothesized to be constructed per word correlates significantly with both reading times and neural activity patterns (Kajikawa et al., 2024; Brennan et al., 2012).
Given these findings, if attention can serve as a general algorithm for memory retrieval in human sentence processing, human memory retrieval should be captured by the attention mechanism operating on syntactic structures as well as that operating on token sequences. In this paper, we investigate whether the attention mechanism of Transformer Grammar (TG; Sartran et al., 2022), which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval, using Normalized Attention Entropy (NAE; Oh and Schuler, 2022) as the linking hypothesis between models and humans. Our experiments demonstrate that TG’s attention achieves superior predictive power for self-paced reading times compared to vanilla Transformer’s, with further analyses revealing independent contributions from both models. These findings suggest that human sentence processing involves dual memory representations—one based on syntactic structures and another on token sequences—with attention serving as the general memory retrieval algorithm, while highlighting the importance of incorporating syntactic structures as representational units.111Code for reproducing our results is available at https://github.com/osekilab/TG-NAE.
2 Background
2.1 Cue-based retrieval
[…]
The worker was surprised that the resident[subj,anim] [who was living near the dangerous warehouse] was complaining about the investigation.
[…]
2.2 Normalized Attention Entropy (NAE)
In recent computational psycholinguistics, attempts have been made to interpret the attention mechanism—a weighted reference of preceding tokens based on Query and Key vectors—as a computational implementation of cue-based retrieval. Notably, Ryu and Lewis (2021) proposed Attention Entropy (AE) as a linking hypothesis, where the diffuseness of attention weights is assumed to quantify the degree of retrieval interference. While AE was initially proposed for modeling interference effects in specific constructions, Oh and Schuler (2022) extended it to naturally occurring text by introducing two normalizations: (i) division by the maximum entropy achievable given the number of preceding tokens, and (ii) sum-to-1 renormalization of attention weights over preceding tokens (Normalized AE, NAE).222Oh and Schuler (2022) showed that regression models for predicting reading times fail to converge with vanilla AE.
where is the set of preceding token positions,333For vanilla Transformers, . is the renormalized attention weight, and represents the attention weight from the -th token (Query) to the -th preceding token (Key) in the -th head of layer .444Oh and Schuler (2022) explored NAE calculation using various attention weight formulations, but in this paper, we adopt the norm-based attention weight formulation (Kobayashi et al., 2020), which achieved the highest predictive power on the self-paced reading time corpus. In this paper, we employ this NAE as a linking hypothesis between attention mechanisms and human memory retrieval.555While Oh and Schuler (2022) also proposed other metrics based on distances between attention weights at consecutive time steps, we exclusively adopt NAE because (i) in TG, the number of preceding elements varies with time, making distance definition non-trivial, and (ii) Oh and Schuler (2022) demonstrated that NAE’s predictive power subsumes that of distance-based metrics in the self-paced reading time corpus.
2.3 Transformer Grammar (TG)
Transformer Grammar (TG; Sartran et al., 2022) is a type of syntactic LM, a generative model that jointly generates token sequences and their corresponding syntactic structures . TG formulates the generation of as modeling a sequence of actions, (e.g., (S (NP The blue bird NP) (VP sings VP) S)), constructing both token sequences and their syntactic structures in a top-down, left-to-right manner. The action sequence comprises three types of operations:
(X: Generate a non-terminal symbol (X, where X represents a phrasal tag such as NP;
w: Generate a terminal symbol w, where w represents a token such as bird;
X): Generate X) to close the most recent opened non-terminal symbol, where X matches the phrasal tag of the targeted non-terminal symbol.
The probability of action sequence is decomposed using the chain rule. Formally, TG is defined as:
TG’s key innovation lies in its handling of closed phrases: immediately after generating X), it computes a vector representation of the closed phrase, which subsequent next-action predictions use as the representation for that phrase. Technically, this operation is realized via two components: X) action duplication and a specialized attention mask. The duplication process transforms into by duplicating all X) actions (e.g., (S (NP The blue bird NP) NP) (VP sings VP) VP) S) S)), while preserving the modeling space by preventing predictions for duplicated positions. The attention mask implements two distinct attention mechanisms: COMPOSE and STACK (Figure 1). COMPOSE operates exclusively at the first occurrence of each X) to generate the phrasal representation by attending only to vectors between the corresponding (X and X) (without making predictions). STACK operates at all other positions to compute representations for next-action prediction, with attention restricted to positions on the stack (comprising unclosed non-terminals, not-composed terminals, and closed phrases).
Previous research has demonstrated that TG’s probability estimates align more closely with human offline grammaticality judgments (Sartran et al., 2022) and online brain activity (Wolfman et al., 2024) than vanilla Transformers. This paper investigates whether the attention mechanism of TG, which uniquely operates on syntactic structures as representational units, can serve as a cognitive model of human memory retrieval.
3 Methods
3.1 NAE calculation with TG
The calculation of NAE with TG requires assumptions regarding two key perspectives:
1.
What syntactic structures should be assumed for a given token sequence?
2.
How should the cognitive load of attention from non-lexical symbols (i.e., (X and X)) be attributed to lexical tokens?
In response to these considerations, we make the following assumptions:
1-A.
We assume only the globally correct syntactic structure (i.e., “perfect oracle”; Brennan, 2016).
2-A.
We consider only attention from lexical tokens, excluding attention from non-lexical symbols.
The adoption of 1-A. is motivated by two factors. First, the self-paced reading time corpus we utilized here provides gold-standard syntactic structures for each sentence, and previous studies have developed predictors based on these annotations (Shain et al., 2020; Isono, 2024). Using the same structural assumptions enables fair comparison with these established predictors, considering the possibility of parsing errors. Second, TG’s current implementation lacks beam search procedure (Stern et al., 2017; Crabbé et al., 2019), an inference technique commonly used in cognitive modeling to handle local ambiguities through parallel parsing (Hale et al., 2018; Sugimoto et al., 2024).666As a proof of concept, we also conducted experiments using multiple syntactic structures generated by word-synchronous beam search with Recurrent Neural Network Grammar (Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021), obtaining similar results (Appendix D).
3.2 Settings
Language models
We used 16-layer, 8-head TG and Transformer (252M parameters).777https://github.com/google-deepmind/transformer_grammars All hyperparameters followed the default settings described in Sartran et al. (2022) (see Appendix A). Following Oh and Schuler (2022), we computed NAE separately for each attention head at the top layer and then summed the values across heads.
Training data
We used BLLIP-lg, a dataset containing 42M tokens (1.8M sentences) from the Brown Laboratory for Linguistic Information Processing (BLLIP) 1987–89 WSJ Corpus Release 1 (Charniak et al., 2000).888https://catalog.ldc.upenn.edu/LDC2000T43 The corpus was re-parsed using a state-of-the-art constituency parser (Kitaev and Klein, 2018) and split into train-val-test sets by Hu et al. (2020). BLLIP-lg has been widely used for training syntactic LMs, including TG. Following Sartran et al. (2022), we trained a 32K SentencePiece tokenizer (Kudo and Richardson, 2018) on the training set and segmented each sentence into subword units.
Both TG and Transformer were trained at the sentence level: TG maximized the joint probability on action sequences, while Transformer maximized the probability on terminal subword sequences. For training hyperparameters, we largely followed the default settings in Sartran et al. (2022) but adjusted the batch size to fit within the memory constraints of our hardware (NVIDIA A100, 40GB). Accordingly, we tuned other hyperparameters (e.g., learning rate) to maintain training stability.999For the detailed hyperparameters, see Appendix A. We trained three models with different random initialization seeds and selected the checkpoint with the lowest validation loss for each run.
Model | LogLik () | Predictor | Effect size [] | -value range | Significant seeds |
TG | 76.6 (8.1) | tg_nae | 1.42 ( 0.2) | 0.001 | 3/3 |
tg_nae_so | 2.26 ( 0.1) | 0.001 | 3/3 | ||
Transformer | 42.8 (9.5) | tf_nae | 1.32 ( 0.2) | 0.001 | 3/3 |
tf_nae_so | 1.46 ( 0.2) | 0.001 | 3/3 |
Reading time data
We used the Natural Stories corpus (Futrell et al., 2018),101010https://github.com/languageMIT/naturalstories. We used the corrected version that addresses the data misalignment issue identified in May 2025. “a series of English narrative texts designed to contain many low-frequency and psycholinguistically interesting syntactic constructions while still sounding fluent and coherent.” We selected this corpus for these “interesting” syntactic constructions, which provide an ideal testbed for investigating memory retrieval effects that might be less pronounced in simpler, more naturalistic texts. The corpus has also been used in several studies investigating memory-related processing mechanisms (Shain et al., 2016; Dotlačil, 2021; Isono, 2024).
The Natural Stories corpus consists of 10 stories (485 sentences, 10,245 words) with self-paced reading times collected from 181 anonymized native English speakers. Following Futrell et al.’s preprocessing, data points were removed if (i) a participant scored less than 5/6 on comprehension questions for a story or (ii) individual reading times were less than 100 or greater than 3,000 . Following Oh and Schuler (2022), we also excluded sentence-initial and sentence-final data points. We further removed sentence-second data points, as they lack the log trigram frequency of the previous token required for our baseline regression model. After preprocessing, 725,875 data points from 180 participants remained for statistical analysis, out of the original 848,875 data points.
Statistical analysis
We evaluated each LM’s NAE contribution to reading time prediction by measuring improvements in regression model fit when adding NAE as predictors. For each model (TG/Transformer), we included both the current word’s NAE (tg_nae/tf_nae) and the previous word’s NAE (tg_nae_so/tf_nae_so) to account for spillover effects (Mitchell, 1984).111111_so indicates spillover.121212Following Oh and Schuler (2022), we summed the subword NAE values for each word. Model improvement was quantified as the increase in log-likelihood (LogLik). This evaluation was conducted for each random seed, and we report the mean LogLik with standard deviation.
Following previous studies such as Dotlačil (2021), Shain et al. (2016), and Isono (2024), the baseline regression model controlled for non-structural, basic aspects of text known to affect reading times:
zone and position (integer): word position in the story and sentence;
wordlen (integer): number of characters in the word;
unigram, bigram, and trigram (continuous): log-transformed n-gram frequencies.
We additionally included the following predictors:
tg_surp and tf_surp (continuous): surprisal from TG and Transformer;
stack_count (integer): number of elements in the stack (comprising unclosed non-terminals, not-composed terminals, and closed phrases).
Following Oh and Schuler (2022), we included surprisal to test NAE’s significance in the presence of surprisal predictors from the same LMs.131313For an experiment on the predictive power of surprisal itself, see Appendix E. Stack count was included to isolate the cost of holding elements Joshi (1990); Abney and Johnson (1991); Resnik (1992) from their interference effects, which TG’s NAE was designed to capture. For the correlations between the predictors, see Appendix B.
All predictors were z-transformed, and we also included the previous word’s values as predictors to model spillover, except for the positional information. The baseline regression model was a linear mixed-effects model (Baayen et al., 2008) with these fixed effects and by-subject and by-story random intercepts:
To assess each LM’s independent contribution to reading time prediction, we also conducted likelihood ratio tests (Wurm and Fisicaro, 2014) by extending Equation 3.2 in two ways: adding both LMs’ NAE versus adding only one LM’s NAE. Note that a larger LogLik from one LM does not necessarily indicate that it contributes above and beyond the other LM, nor does a smaller LogLik indicate no unique contribution. Following Aurnhammer and Frank (2019), we used NAE and surprisal values averaged across random seeds for these nested model comparisons.
4 Results
4.1 Does TG’s NAE have predictive power for reading times?
Table 1 presents the contributions of TG’s and Transformer’s NAE to reading time prediction. First, Transformer’s NAE exhibited significant predictive power for reading times, independent of baseline predictors such as surprisal. The effect size was in the expected positive direction (higher NAE values corresponding to longer reading times), showing both immediate and spillover effects. This corroborates the arguments of Ryu and Lewis (2021) and Oh and Schuler (2022) that the attention mechanism—the weighted reference of preceding tokens—functions as a cognitive model of human memory retrieval, despite its engineering-oriented origins.
Second, TG’s NAE exhibited robust predictive power, demonstrating significant positive effects in both immediate and spillover conditions. This finding not only provides additional evidence for incremental construction of syntactic structures in human sentence processing (e.g., Fossum and Levy, 2012), but also suggests that TG’s attention mechanism effectively models memory retrieval from these constructed syntactic representations.
Finally, TG’s NAE made a substantially stronger contribution to reading time prediction (LogLik76.6) compared to Transformer’s NAE (LogLik42.8). This finding suggests that retrieval from syntactic memory representations plays a more dominant role in human sentence processing than retrieval from lexical memory representations. This underscores the importance of incorporating syntactic structures as a unit of memory representation, which we implemented through the integration of TG and NAE here.
4.2 Do TG’s and Transformer’s NAE have independent contributions?
Figure 2 presents the results of likelihood ratio tests examining the independence of TG’s and Transformer’s NAE contributions. The regression model incorporating NAE from both LMs (‘TG & Transformer’) demonstrated significantly higher predictive power than the models containing NAE from either LM alone (‘TG’ or ‘Transformer’). This reveals that TG’s NAE certainly captures variance in reading times that Transformer’s NAE cannot explain, while Transformer’s NAE, despite its lower overall predictive power, accounts for unique variance not captured by TG’s NAE. This finding aligns with psycholinguistic literature, where cognitive models of memory retrieval encompass both syntax-based approaches (e.g., verb-argument relationships; Lewis and Vasishth, 2005) and semantic-based approaches (e.g., bag-of-words-like similarity; Brouwer et al., 2012), suggesting that the attention mechanisms of TG and Transformer serve as complementary cognitive models, each capturing distinct aspects of human memory retrieval.
4.3 What aspects of memory retrieval do TG’s and Transformer’s NAE capture?
To investigate the aspects of human memory retrieval captured by TG’s and Transformer’s NAE, we analyzed differences in prediction improvement across part-of-speech (POS) tags annotated in the Natural Stories corpus.141414For a complete list of POS tags in the Natural Stories corpus, see Appendix C. Our analysis followed three steps: (i) selecting POS tags with more than 1,000 occurrences, (ii) for each POS tag, testing the significance of improvement from the baseline regression model (measured in Root Mean Squared Error, RMSE) when adding NAE of the current and previous word as fixed effects,151515We used the same regression models as in Section 4.2, where surprisal and NAE values were averaged across seeds. and (iii) examining the significance of differences in RMSE between TG and Transformer for POS tags where either model showed significant improvement. We assessed significance using Wilcoxon signed-rank tests with Bonferroni correction ().
Figure 3 presents the differences in prediction improvement across POS tags. Consistent with the larger LogLik value, TG’s NAE demonstrated advantages over Transformer’s NAE across a broader range of POS tags. Notably, TG’s NAE exhibited superior improvement across verbs (VB, VBG, VBN, and VBP), while Transformer’s NAE excelled for nouns (NN and NNP). This pattern aligns with our earlier argument regarding the complementary nature of these models (Section 4.2), indicating that different types of retrieval operations—verb-triggered retrieval, which often relies on syntactic features (e.g., argument structure), and noun-triggered retrieval, which often relies on semantic features (e.g., referential associations)—are better captured by distinct attention mechanisms: attention with syntactic and token memory representations, respectively.
5 Follow-up analysis
5.1 Do TG’s advantages stem from the COMPOSE attention?
Model | LogLik | Predictor | -value |
TG | 46.1 | *_nae | ∗∗ (2/3) |
( 9.1) | *_nae_so | ∗∗∗ (3/3) | |
TG | 18.1 | *_nae | ∗∗ (1/3) |
( 9.3) | *_nae_so | ∗∗∗ (3/3) |
As described in Section 2.3, TG’s key feature is the COMPOSE attention, which explicitly generates single vector representations for closed phrases. Here, we investigate whether TG’s predictive power derives from merely considering syntactic structures or from explicitly treating closed phrases as single representations (see Hale et al., 2018; Brennan et al., 2020). To address this question, we developed TG, a TG variant that processes each action in the action sequence as an individual token without the COMPOSE attention (i.e., Choe and Charniak’s Parsing as Language Modeling approach). We trained TG with identical hyperparameters as TG. The baseline regression model (Equation 3.2) was augmented with (i) TG’s surprisal and (ii) Transformer’s NAE to (i) ensure a fair comparison between TG and TG and (ii) distinguish between the effects of direct terminal token access and syntactic structure consideration in TG.161616For direct comparison between TG and TG under the baseline regression model without Transformer’s NAE, see Appendix F.
Table 2 presents the LogLik values obtained when incorporating either TG’s or TG’s NAE as fixed effects into the baseline regression model. Note that due to the potential multicollinearity between Transformer’s NAE and TG/TG’s NAE, we focus on the LogLik values and significance of the contribution rather than individual effect sizes.
Our analysis reveals two key findings. First, TG’s NAE demonstrates significant predictive power for reading times, even in the presence of Transformer’s NAE, implying that consideration of syntactic structures alone captures certain memory retrievals that the token-level attention mechanism cannot capture. Second, TG’s NAE outperforms TG’s, suggesting that the attention mechanism that treats closed phrases as single representations more effectively captures variance in syntax-based memory retrieval. The likelihood ratio tests further revealed that TG’s NAE captured reading time patterns unexplainable by TG (‘TG & TG’‘TG’, ), while TG did not explain unique variance beyond what TG already accounts for (‘TG & TG’‘TG’, ).
We analyzed the RMSE differences across POS tags to investigate which aspects of human memory retrieval are better captured by the COMPOSE attention. Figure 4 showed that TG’s NAE was consistently superior to TG’s NAE across verbs (VB, VBG, and VBN), highlighting the critical role of the COMPOSE attention in capturing verb-triggered retrieval, a type of retrieval that was identified as distinctively better captured by TG compared to vanilla Transformers (Section 4.3).
5.2 Does TG’s NAE capture interference effects?
Psycholinguistic research has considered two primary types of memory retrieval costs: interference effects, which NAE aims to capture, and decay effects—the cognitive load associated with accessing elements at greater linear distances (e.g., Gibson, 1998, 2000). Here, we examine whether TG’s NAE genuinely captures interference effects by testing its independence from variables that model memory decay effects. For modeling decay effects, we employed Category Locality Theory (CLT; Isono, 2024),171717Although Dependency Locality Theory (DLT; Gibson, 1998, 2000) is widely recognized as one of the most prominent models for capturing decay effects, we opted for CLT here, following Isono’s finding that DLT-based predictors fail to achieve statistical significance in explaining reading times in the Natural Stories corpus. which treats phrases in syntactic structure181818CLT assumes syntactic structure based on Combinatory Categorial Grammar (Steedman, 2000). as representational units of memory and quantifies decay effects using the distance (measured in content words) between an input and the phrases to be composed with it.
To assess independence, we tested whether TG’s NAE and CLT maintain their contributions when simultaneously included in the baseline regression model (Equation 3.2), and examined their independence through likelihood ratio tests.191919As in other likelihood ratio tests, we used surprisal and NAE values averaged across random seeds. The results (Table 3) show that TG’s NAE exhibited significant effects in both immediate and spillover conditions, and CLT demonstrated a significant spillover effect. The likelihood ratio tests confirmed that these effects were independent (‘TG & CLT’‘CLT’, ; ‘TG & CLT’‘TG’, ).
These results provide empirical evidence that NAE quantifies interference rather than decay in memory retrieval—extending beyond previous studies on NAE (Ryu and Lewis, 2021; Oh and Schuler, 2022). This finding is significant because, as far as we are aware, while psycholinguistics has developed various implementations of memory decay effects, it has lacked broad-coverage implementations of interference effects applicable to naturally occurring texts. Our results suggest that NAE represents a promising approach for quantifying interference effects in a broad-coverage manner.
Model | Predictor | Effect size [] |
TG & CLT | tg_nae | ∗∗∗ |
tg_nae_so | ∗∗∗ | |
clt | ||
clt_so | ∗∗∗ |
6 Level of description
In cognitive modeling studies based on surprisal theory, explanations typically follow the form “if these LMs were models of human prediction, the difficulty of next-word disambiguation that humans solve would be approximated as follows.” Such explanations typically operate at the most abstract of Marr’s three levels of description—the computational level. Recently, Futrell et al. (2020) proposed lossy-context surprisal to integrate memory representation perspectives into surprisal theory. However, as they explicitly stated, this theory remains at the computational level, relaxing assumptions about memory representations in human predictive processing. In contrast, cognitive models of human memory, such as cue-based retrieval, generally provide explanations about mechanisms that deal with specific mental representations. These explanations typically move down one level to the algorithmic level of description.
While not explicitly stated by the authors, we argue that the work of Ryu and Lewis (2021) and Oh and Schuler (2022)—conceptualizing attention mechanisms as implementations of cue-based retrieval—represents movement toward the algorithmic level, approaching cue-based retrieval itself. Our research advances this direction by investigating a fundamental question at this level: the nature of memory representations (see Hale, 2014). Future work could fully operate at the algorithmic level by incorporating more realistic elements such as parallel parsing, left-corner parsing strategies, and memory decay mechanisms.
7 Conclusion
In this paper, we have demonstrated that attention can serve as the general algorithm for modeling human memory retrieval from two representational systems. Furthermore, we have shown that among the LMs examined in this paper (TG, TG, and Transformer), TG—whose attention mechanism uniquely operates on syntactic structures as representational units—best captures dominant factors in human sentence processing. Our results suggest that the integration of attention mechanisms (developed in NLP) with syntactic structures (theorized in linguistics) constitutes a broad-coverage candidate implementation for human memory retrieval. We hope these findings will foster greater collaboration between these two fields.
Limitations
Our NAE calculation comprised three steps: (i) computing NAE for each attention head in the top layer, (ii) adding the values across heads, and (iii) summing subword-level values into word level. While this procedure strictly adhered to Oh and Schuler (2022), alternative approaches to handling layers, attention heads (Ryu and Lewis, 2021), and subword tokens (Oh and Schuler, 2024; Giulianelli et al., 2024) warrant investigation.
While our study provides an in-depth investigation using the Natural Stories corpus—an English self-paced reading time corpus containing many interesting syntactic constructions—the breadth of our analysis has certain limitations. The generalizability of our findings to different languages (e.g., Japanese self-paced reading time corpus from Asahara, 2022) and other cognitive load (e.g., gaze duration from Kennedy et al., 2003 or EEG and fMRI from Bhattasali et al., 2020) remains to be investigated.
As discussed in Section 3.1, we employed “perfect oracles” as syntactic structures behind token sequences. While this assumption has been widely adopted in previous studies on human syntactic processing (Brennan, 2016; Shain et al., 2016; Stanojević et al., 2021; Isono, 2024), this idealization leaves the resolution of local ambiguities, which humans encounter during actual sentence processing, outside the scope of our study (for a conceptual case study, see Appendix D).
Although we adopted the default TG implementation of a top-down parsing strategy, psycholinguistic literature has suggested that a left-corner parsing strategy might be more plausible for human sentence processing from a perspective of memory capacity (cf. stack_count) (Abney and Johnson, 1991; Resnik, 1992). However, when contrasting memory representations based on syntactic structures versus token sequences, the attention mechanism of the top-down TG can serve as a reasonable approximation of retrieval from structure-based memory—this aligns with previous work contrasting predictions based on syntactic structures versus token sequences, which used top-down TG or RNNG to represent structure-based prediction (Wolfman et al., 2024; Hale et al., 2018; Brennan et al., 2020).
Finally, while this paper focused on investigating the attention mechanism through the lens of memory-based theory, exploring TG and Transformer as integrated implementations for expectation-based theory (via surprisal) and memory-based theory (via NAE) represents a promising future direction (Michaelov et al., 2021; Ryu and Lewis, 2022). Specifically, future work could investigate the attention mechanism as the underlying driver of surprisal’s predictive power (Appendix E), analyzing the relationship between surprisal and NAE.
Ethical considerations
We employed AI-based tools (Claude, ChatGPT, GitHub Copilot, and Grammarly) for writing and coding assistance. These tools were used in compliance with the ACL Policy on the Use of AI Writing Assistance.
Acknowledgments
We sincerely thank Michael Wolfman and John Hale for providing detailed information on Wolfman et al. (2024). We also appreciate the insightful reviews provided by anonymous ARR 2025 February reviewers. This work was supported by JSPS KAKENHI Grant Number 24H00087, Grant-in-Aid for JSPS Fellows JP24KJ0800, JST PRESTO Grant Number JPMJPR21C2, and JST SPRING Grant Number JPMJSP2108.
Appendix A Hyperparameters
Model architecture | Transformer-XL (Dai et al., 2019) |
Vocabulary size | 32,768 |
Model dimension | 1,024 |
Feed-forward dimension | 4,096 |
Number of layers | 16 |
Number of heads | 8 |
Segment length | 256 |
Memory length | 256 |
Optimizer | Adam (, ) (Kingma and Ba, 2015) |
Batch size | 16 |
Number of training steps | 400,000 |
Learning rate scheduler | Linear warm-up & cosine annealing |
Number of warm-up steps | 32,000 |
Initial learning rate | |
Maximum learning rate | |
Final learning rate | |
Dropout rate | 0.1 |
The hyperparameters are shown in Table 4. All model hyperparameters follow Sartran et al. (2022); Wolfman et al. (2024), while training hyperparameters were adjusted to accommodate the batch size suitable for our computational resources (NVIDIA A100, 40GB). The total computational cost required for all experiments was approximately 225 GPU hours.
Appendix B Correlations between the predictors
The correlations between the predictors in our statistical analysis are shown in Table 5. While the NAE from different LMs shows very high correlations with each other, TG and Transformer provide independent predictive power for the self-paced reading times; TG subsumes the predictive power of TG (see Section 4.2 and 5.1).202020_mcomp indicates .
Appendix C Part-of-speech tags
Table 5 presents the complete list of part-of-speech (POS) and symbol tags in the Natural Stories corpus. As reading times are annotated for each whitespace-delimited region, for data points containing symbol tags (e.g., NNP.), we used the stripped version (e.g., NNP) in our analysis. Additionally, we excluded from our analysis any data points containing multiple POS tags (e.g., NNP POS).
CC | Coordinating conjunction | PRP$ | Possessive pronoun |
CD | Cardinal number | RB | Adverb |
DT | Determiner | RBR | Adverb, comparative |
EX | Existential there | RBS | Adverb, superlative |
FW | Foreign word | RP | Particle |
IN | Preposition or subordinating conjunction | TO | to |
JJ | Adjective | UH | Interjection |
JJR | Adjective, comparative | VB | Verb, base form |
JJS | Adjective, superlative | VBD | Verb, past tense |
MD | Modal | VBG | Verb, gerund or present participle |
NN | Noun, singular or mass | VBN | Verb, past participle |
NNS | Noun, plural | VBP | Verb, non-3rd person singular present |
NNP | Proper noun, singuler | VBZ | Verb, 3rd person singular present |
NNPS | Proper noun, plural | WDT | Wh-determiner |
PDT | Predeterminer | WP | Wh-pronoun |
POS | Possessive ending | WP$ | Possessive wh-pronoun |
PRP | Personal pronoun | WRB | Wh-adverb |
-LRB- | Left round bracket | , | Comma |
-RRB- | Right round bracket | . | Period |
‘‘ | Open double quotes | : | Colon |
’’ | Closing double quotes |
Appendix D Parallel parsing experiment
Model | LogLik () | Predictor | Effect size [] | -value range | Significant seeds |
TG | 56.0 (5.5) | tg_bs_nae | 1.08 ( 0.1) | 0.001 | 3/3 |
tg_bs_nae_so | 2.01 ( 0.1) | 0.001 | 3/3 | ||
Transformer | 26.1 (8.1) | tf_nae | 0.95 ( 0.1) | 0.001 | 3/3 |
tf_nae_so | 1.25 ( 0.2) | 0.001 | 3/3 |
As a conceptual case study for the local ambiguity resolution in syntactic structures behind token sequences, we implemented TG’s NAE calculation using 10 syntactic structures obtained through word-synchronous beam search (Stern et al., 2017) with Recurrent Neural Network Grammar (RNNG; Dyer et al., 2016; Kuncoro et al., 2017; Noji and Oseki, 2021).212121https://github.com/aistairc/rnng-pytorch222222RNNG was trained on BLLIP-lg using default hyperparameters. For inference, action beam size and fast track were set to 100 and 1, respectively. NAE was computed individually for each syntactic structure and then aggregated as a weighted average:
where represents the set of syntactic structures synchronized at the -th word (), and denotes the token position corresponding to the -th word in structure .232323stack_count was similarly calculated as the weighted average across syntactic structures in .
The analysis revealed patterns consistent with those observed when considering only the globally correct syntactic structure: both LMs’ NAE demonstrated significant predictive power for reading times, with TG’s NAE showing stronger contributions compared to Transformer’s (Table 6).242424_bs indicates beam search. The likelihood ratio tests further confirmed independent contributions from both LMs ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’).
Appendix E Surprisal experiment
Model | LogLik () | Predictor | Effect size [] | -value range | Significant seeds |
TG | 62.6 (6.8) | tg_surp | 1.36 ( 0.0) | 0.001 | 3/3 |
tg_surp_so | 1.88 ( 0.1) | 0.001 | 3/3 | ||
Transformer | 159 (12) | tf_surp | 2.47 ( 0.1) | 0.001 | 3/3 |
tf_surp_so | 2.98 ( 0.2) | 0.001 | 3/3 |
We analyzed each LM’s surprisal contribution to reading time prediction using a baseline regression model that excluded both LMs’ surprisal from Equation 3.2 but included their NAE (Table 7). While both LMs’ surprisal demonstrated significant predictive power for reading times, Transformer’s surprisal exhibited a stronger contribution compared to TG’s. Additionally, our likelihood ratio tests using the averaged surprisal revealed that the regression model incorporating both LMs’ surprisal showed significantly higher predictive power compared to models with either LM’s surprisal alone ( for both comparisons: ‘TG & Transformer’‘Transformer’ and ‘TG & Transformer’‘TG’). These findings suggest that (i) unlike attention mechanisms, next-token prediction based solely on token sequences more effectively captures dominant factors of human predictive processing, but (ii) similar to attention mechanisms, both types of next-token prediction—those based on token sequences alone and those leveraging both syntactic structures and token sequences—may coexist as models that capture distinct aspects of human predictive processing.
Appendix F Direct comparison of TG and TG
Model | LogLik () | Predictor | Effect size [] | -value range | Significant seeds |
TG | 78.2 (6.9) | tg_nae | 1.43 ( 0.2) | 0.001 | 3/3 |
tg_nae_so | 2.28 ( 0.3) | 0.001 | 3/3 | ||
TG | 59.4 (17) | tg_mcomp_nae | 1.36 ( 0.2) | 0.001 | 3/3 |
tg_mcomp_nae_so | 1.90 ( 0.2) | 0.001 | 3/3 |
In Section 5.1, we explored the advantage of treating closed phrases as single representations beyond syntactic structure consideration. Our analysis incorporated Transformer’s NAE in the baseline regression model to distinguish between two effects in TG: syntactic structure consideration and direct terminal token access.
To evaluate which model—TG or TG—better captures more dominant factors in human sentence processing as a single model, we assessed their predictive power without Transformer’s NAE in the baseline regression model (Table 8). The analysis revealed TG’s superior predictive power (LogLik=78.2) compared to TG (LogLik=59.4). These results highlight that TG, which explicitly treats closed phrases as single representations, outperforms TG, even when considering TG’s advantage in direct terminal token access. The likelihood ratio tests confirmed TG’s independent predictive power from TG (‘TG & TG’‘TG’, ), while in contrast to Section 5.1, TG also accounted for unique variance (‘TG & TG’‘TG’, ). This bidirectional relationship likely emerges because TG’s direct terminal token access explains unique variance in the absence of Transformer’s NAE.
Appendix G License
Dataset/Tool | License |
BLLIP (Charniak et al., 2000) | BLLIP 1987–89 WSJ Corpus Release 1 |
Natural Stories corpus (Futrell et al., 2018) | CC BY-NC-SA 4.0 |
transformer_grammar (Sartran et al., 2022) | Apache 2.0 |
rnng-pytorch (Noji and Oseki, 2021) | MIT License |
SentencePiece (Kudo and Richardson, 2018) | Apache 2.0 |
R (version 4.4.2) (R Core Team, 2024) | GNU GPL 2 |
lme4 (version 1.1.34) (Bates et al., 2015) | GNU GPL 2 |
lmerTest (version 3.1.3) (Kuznetsova et al., 2017) | GNU GPL 2 |
Table 9 summarizes the licenses of the data and tools employed in this paper. All data and tools were used under their respective license terms.